two side of a triangle are √3 - 1 and √3 + 1 units and their included angle is 60° . solve the triangle . ?

Grade:11

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Arunav

Cos60=(√3 +1)2 + (√3 -1)2 -a2/2(√3 +1)(√3 -1)

solving a=√6

now use Cos formula to find out other angles like this

 

 All the best.                                                          

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

 

vikas askiitian expert
13 years ago

let a,b are 31/2-1 , 31/3+1 & C = 60o        then 

cosC = a2+b2-c2/2ab

ab = a2+b2-c2

3-1 + 6+2 -c2

c2 = 6

c = 61/2..................................1

now using a/sinA = b/sinB = c/sinC =k

     sinA = asinC/c = (31/2-1/2*21/2)

       A = 15o                ..............................2

    A+B+C = 180

    B + 60+15 = 180

   B = 105o..........................3